# Multimodal Fact-Checking: An Agent-based Approach

### v2

@@ I Introduction
The proliferation of social media has intensified the spread of persuasive textual and visual misinformation, posing serious risks to public health and behavior [borges2022infodemics, wsj2025falseTariffHeadline, ARCURI2023106130]. Advances in generative AI further exacerbate this threat by enabling realistic, multimodal fake content [berman2024aimisinformation, DeepfakeSurvey2024Wang, reuters2025duterteDisinformation]. Traditional binary fact-checking is inadequate, as it lacks rationales and evidence that foster public trust [blanchar2024trump, Porter2021]. Narrow tasks like out-of-context (OOC) detection [qi2024sniffer] and deepfake detection [NoiseDF2023Wang, wang2023dire] fall short in addressing real-world complexities. To enhance transparency and reliability, explainable multimodal fact-checking has become a crucial research focus.
Recent work has explored LLM-based methods using techniques such as chain-of-thought prompting, claim decomposition, question-guided prompting, and knowledge distillation [PanQACheck23, leite2023detecting, zhang2023towards, kareem2023fighting]. LVLM-based frameworks have extended these techniques to multimodal settings by incorporating image search and external retrieval, evolving into agent-collaborative systems [wang2024mmidr, xuan2024lemma, tonglet2024image, qi2024sniffer, DuICML24, braun2025defame].
@@ II-B Misinformation and OOC detection methods
Approaches to misinformation and out-of-context (OOC) detection can be classified along two axes: (1) internal estimation vs. external retrieval, and (2) feature-based vs. LLM/LVLM-based reasoning. Feature-based internal methods rely on writing style, propagation patterns, or image-text consistency [przybyla2020capturing, shu2020hierarchical, zhou2020multimodal, luo2021newsclippings], but lack external grounding and robustness. Feature-based external approaches retrieve related content (e.g., captions, images) and compare via semantic or entity-level similarity [aneja2023cosmos, muller2020multimodal, Papadopoulos_2025_WACV, abdelnabi2022open], though often lacking deeper reasoning. LLM/LVLM-based internal reasoning uses prompting and knowledge distillation [PanQACheck23, leite2023detecting, zhang2023towards], but is limited by internal knowledge and modality coverage. LLM/LVLM-based external methods leverage instruction-tuned models or agent-based designs with retrieval [qi2024sniffer, DuICML24, xuan2024lemma, braun2025defame], enhancing interpretability but still falling short on multi-step evidence synthesis.
@@ IV-B Agents for Fact-Checking
As illustrated in Figure 4, the framework consists of five specialized agents: 1) Strategy Planning Agent (Agent-SP) , 2) Text Evidence Retrieval and Validation Agent (Agent-TR), 3) Image Retrieval and Analysis Agent (Agent-IR), 4) Reasoning Agent (Agent-R), 5) Explanation Generation Agent (Agent-EG). Each agent is designed with task-specific knowledge to mimic distinct components of real-world fact-checking workflows, and is deployed using prompt-engineering techniques with tool interaction. The simplified prompts are provided in the Appendix VI.
@@ V-B Implementation Details
Data Leakage Prevention: To ensure fair evaluation and prevent data leakage, we apply strict filtering to all retrieved evidence used before final reasoning for all external evidence based methods. Specifically, we exclude: (1) any evidence retrieved from Snopes or other major fact-checking websites (see details in Appendix VI), and (2) any evidence published after the original fact-checking date of the claim.
@@ V-D Performance on Explainability
so that higher values consistently indicate better performance. Full annotation guidelines and the interface are provided in the Appendix VI.
@@ V-E Ablation Study of Agents
We provide additional analysis on model performance across varying levels of sample difficulty in the Appendix VI.
@@ VI Conclusion
First, as discussed ine0989389@u.nus.edu the case study section, current agent-based fact-checking methods still fall short in understanding real-world contextual nuances of claims and evidence, which is crucial for more intelligent and reliable fact-checking.
Second, processing multiple pieces of evidence and images incurs considerable computational cost, which limits scalability. Exploring open-source models and lightweight deployment strategies is practically beneficial for reducing costs and supporting accessible fact-checking.

### v3

@@ I Introduction
The proliferation of social media has intensified the spread of persuasive textual and visual misinformation, posing serious risks to public health and behavior [borges2022infodemics, wsj2025falseTariffHeadline, ARCURI2023106130]. Advances in generative AI further exacerbate this threat by enabling realistic, multimodal fake content [berman2024aimisinformation, DeepfakeSurvey2024Wang, reuters2025duterteDisinformation, 11010889]. Traditional binary fact-checking is inadequate, as it lacks rationales and evidence that foster public trust [blanchar2024trump, Porter2021]. Narrow tasks like out-of-context (OOC) detection [qi2024sniffer] and deepfake detection [NoiseDF2023Wang, 10487975, 10290956, 9408664] fall short in addressing real-world complexities. To enhance transparency and reliability, explainable multimodal fact-checking has become a crucial research focus.
Recent work has explored LLM-based methods using techniques such as chain-of-thought prompting, claim decomposition, question-guided prompting, and knowledge distillation [PanQACheck23, leite2023detecting, zhang2023towards, kareem2023fighting]. LVLM-based frameworks have extended these techniques to multimodal settings by incorporating image search and external retrieval, evolving into agent-collaborative systems [wang2024mmidr, xuan2024lemma, 10925399, tonglet2024image, qi2024sniffer, DuICML24, braun2025defame].
@@ II-B Misinformation and OOC detection methods
Approaches to misinformation and out-of-context (OOC) detection can be classified along two axes: (1) internal estimation vs. external retrieval, and (2) feature-based vs. LLM/LVLM-based reasoning. Feature-based internal methods rely on writing style, propagation patterns, or image-text consistency [przybyla2020capturing, shu2020hierarchical, zhou2020multimodal, luo2021newsclippings, 9996587], but lack external grounding and robustness. Feature-based external approaches retrieve related content (e.g., captions, images) and compare via semantic or entity-level similarity [aneja2023cosmos, muller2020multimodal, Papadopoulos_2025_WACV, abdelnabi2022open], though often lacking deeper reasoning. LLM/LVLM-based internal reasoning uses prompting and knowledge distillation [PanQACheck23, leite2023detecting, zhang2023towards], but is limited by internal knowledge and modality coverage. LLM/LVLM-based external methods leverage instruction-tuned models or agent-based designs with retrieval [qi2024sniffer, DuICML24, xuan2024lemma, braun2025defame], enhancing interpretability but still falling short on multi-step evidence synthesis.
@@ IV-B Agents for Fact-Checking
As illustrated in Figure 4, the framework consists of five specialized agents: 1) Strategy Planning Agent (Agent-SP) , 2) Text Evidence Retrieval and Validation Agent (Agent-TR), 3) Image Retrieval and Analysis Agent (Agent-IR), 4) Reasoning Agent (Agent-R), 5) Explanation Generation Agent (Agent-EG). Each agent is designed with task-specific knowledge to mimic distinct components of real-world fact-checking workflows, and is deployed using prompt-engineering techniques with tool interaction. The simplified prompts are provided in the Appendix Multimodal Fact-Checking: An Agent-based Approach.
@@ V-B Implementation Details
Data Leakage Prevention: To ensure fair evaluation and prevent data leakage, we apply strict filtering to all retrieved evidence used before final reasoning for all external evidence based methods. Specifically, we exclude: (1) any evidence retrieved from Snopes or other major fact-checking websites (see details in Appendix Multimodal Fact-Checking: An Agent-based Approach), and (2) any evidence published after the original fact-checking date of the claim.
@@ V-D Performance on Explainability
so that higher values consistently indicate better performance. Full annotation guidelines and the interface are provided in the Appendix Multimodal Fact-Checking: An Agent-based Approach.
@@ V-E Ablation Study of Agents
We provide additional analysis on model performance across varying levels of sample difficulty in the Appendix Multimodal Fact-Checking: An Agent-based Approach.
@@ VI Conclusion
First, as discussed in the case study section, current agent-based fact-checking methods still fall short in understanding real-world contextual nuances of claims and evidence, which highlights more intelligent and reliable fact-checking.
Second, processing multiple pieces of evidence and images incurs considerable computational cost, which limits scalability. Exploring open-source models and lightweight deployment strategies will be more practically beneficial for fact-checking.